OpenStudy (anonymous):

A student drops a rock from a bridge to the water 12 m below. With what speed does the rock strike the water? Answer in units of m/s

OpenStudy (jamesj):

Well, how much Kinetic Energy does the rock have when it strikes the water?

OpenStudy (anonymous):

conservation energy. K(kinetic energy)=U(potential energy) . i think...

OpenStudy (jamesj):

Yes, so much much gravitational potential energy has been converted into KE. Use that to calculate the rock's speed.

OpenStudy (shane_b):

I'm not sure how conservation of energy helps here since you aren't given the mass. I'd just use the kinematic equation:\[\large V_f^2=V_i^2+2ad\]

OpenStudy (jamesj):

And where, pray tell, does that equation come from? It comes directly from Conservation of Energy!

OpenStudy (shane_b):

As I said, you're not given a mass here. If there's a way to solve it with COE I'm just not seeing :)

OpenStudy (jamesj):

The loss of potential energy of the rock is PE = mgh, where h = 12m The KE the rock has is (1/2)mv^2 This must be equal to the change in PE. Hence \[ mgh = \frac{1}{2}mv^2 \] and therefore \[ v^2 = 2gh \] Now solve for the speed, v

OpenStudy (shane_b):

Yea, as soon as I posted that I realized the masses cancel...dumb me.

OpenStudy (jamesj):

and by the way, this is where that kinematic equation comes from. By the Work-Energy, theorem, the change in kinetic energy of an object is--barring other sorts of energy being involved--equal to the work done on it. The work done is W = Fd = mad, F = force, d = distance, m = mass, a = acceleration The change in KE is (1/2)mv^2 - (1/2)mu^2, where v = final velocity, u = initial velocity Thus as W = change in KE \[ mad = \frac{1}{2}m(v^2 - u^2) \] Canceling mass terms and rearranging, \[ v^2 = u^2 + 2ad \]